• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name P. Monaghan

    Log entry time 06:48:41 on March26,2005

    Entry number 142448

    keyword=Compton not looking good?

    Hachemi came in around 5:45am and rebooted the compton computer for us and set up the controls guis again - he had been up and read my earlier halog about the compton being frozen ( at 5am!! ).

    Unfortunately, I noticed at around 6:30am, shortly after Hachemi left, that the electron and photon rates had flatlined again. I ended the current coda run on compton and restarted a new run. This seemed to bring the readbacks back to life again, but they now look rather ragged. The stripchart and control monitor gui are shown below. Call the counting house (ext. 6328) with suggestions please.


    FIGURE 1

    FIGURE 2